OpenStudy (anonymous):

Find the first four terms of a squence using the recursive definition. f(1)=6, f(n)=f(n-1)-5

OpenStudy (amistre64):

what does the definition tell us?

OpenStudy (anonymous):

Using a process that can be repeated. So like the arithmetic sequence. The arithmetic sequence is adding a constant of the previous term.

OpenStudy (amistre64):

ok, in this case, what are we 'adding' to one term to make the next term? any ideas?

OpenStudy (amistre64):

or if we simply let n={1,2,3,4} and the rule defined, we get: f(1) = 6 f(2) = f(1) -5 f(3) = f(2) -5 f(4) = f(3) -5

OpenStudy (amistre64):

so, what would our 4 terms be?

OpenStudy (anonymous):

So since its asking for the first four terms wouldnt it be f(1)=6 f(n)=f(n-1)-5 f(2)=f(2-1)-5=f(1)+1= 6+1=7 so 7 would be one of the four terms but I dont know how to get the rest.

OpenStudy (amistre64):

f(2) = f(2-1) - 5 f(2) = f(1) - 5 f(2) = 6 - 5 = 1, not 7 so, f(1) = 6, f(2) = 1 ---------------------- now repeat with n=3 f(3) = f(3-1) - 5 f(3) = f(2) - 5 f(3) = 1 - 5 = -4 so, f(1) = 6, f(2) = 1, f(3) = -4 what is f(4) ?

OpenStudy (amistre64):

as we find each new term, we can use it to find the next one .... hence the naming it: 'recursive'

OpenStudy (amistre64):

f(4) = f(4-1) - 5 f(4) = f(3) - 5 ^^^ but we know f(3) = -1 f(4) = -1 - 5

OpenStudy (amistre64):

pfft, f(3) = -4 .... these tired old eyes play tricks on me

OpenStudy (amistre64):

can you tell me how we are working this? its basic to me so i cant really see the difficulty. you have to tell me what it is that is confusing you

OpenStudy (anonymous):

Oka, I understand but im confused on f(4)

OpenStudy (amistre64):

ok, tell me how you are looking at f(4) show me your working

OpenStudy (amistre64):

if you know how to get from f(1) to f(2), then the process does not change. f(2) allows us to get f(3), and f(3) allows us to get f(4). typing errors aside, its pretty repetitive.

OpenStudy (anonymous):

So f(4)=-6?

OpenStudy (amistre64):

f(4) = f(3) - 5 f(3) = -4 [i mistyped it before as -1] f(4) = -4 -5

OpenStudy (amistre64):

another way to look at it, just subtract 5 from the setup before ... f(1) = 6 f(2) = 6-5 f(3) = 6-5-5 f(4) = 6-5-5-5

OpenStudy (anonymous):

So f(4)=-9

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

now, what are our 4 terms?

OpenStudy (anonymous):

1,-5,-4,-9

OpenStudy (amistre64):

f(1) = 6 f(2) = 6-5 = 1 f(3) = 6-5-5 = -4 f(4) = 6-5-5-5 = -9

OpenStudy (anonymous):

oh im sorry I didnt mean to write -5 but thank you so much

OpenStudy (amistre64):

\[\begin{pmatrix}n\\f(n)\end{pmatrix}=\begin{pmatrix}1&2&3&4\\6&1&-4&-5\end{pmatrix}\]

OpenStudy (amistre64):

good luck :)

OpenStudy (amistre64):

yeah, 6, 1 -4, -9 math and typing dont mix that well

OpenStudy (anonymous):

I agree on that ,Do you think you can help me with a couple more?

OpenStudy (amistre64):

maybe one more; its late (1235) and work comes early in the morning

OpenStudy (anonymous):

Okay thank you. f(1)=2,f(n)=-3f(n-1)+[f(n-1)]^2

OpenStudy (amistre64):

well this ones just a exercise in plugging in the value ... show me your work for n=2

OpenStudy (amistre64):

dunno if it helps any to rewrite it by factoring: f(n) = f(n-1) [f(n-1) -3] your call

OpenStudy (anonymous):

Wait okay so do I plug 2 in for n?

OpenStudy (amistre64):

ideally yes, since 2 is the number after 1.

OpenStudy (anonymous):

Okay so I have f(2)=-3f(2-1)+[f(2-1)]^2

OpenStudy (amistre64):

good, and since 2-1 = 1 f(2) = -3 f(1) + [f(1)]^2 and what does f(1) equal?

OpenStudy (anonymous):

Im not sure in what to do after

OpenStudy (amistre64):

you replace f(1) with what its vale is, and work the math

OpenStudy (amistre64):

*value

OpenStudy (anonymous):

So would Imultiply -3 by 1? =-3

OpenStudy (anonymous):

Or where would I get the value from?

OpenStudy (amistre64):

f(1) has been defined for you already in your setup f(1)=2 <------ f(n)=-3f(n-1)+[f(n-1)]^2 ----------------- f(2) = -3 f(1) + [f(1)]^2 f(2) = -3(2) + 2^2 = -6+4 = -2 f(2) = -2 ------------------ we know f(2) now, so let n=3 and work the process again

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (amistre64):

youre welcome .... once we know a new term, we can use it to find the next new term f(3) = -3 f(2) + [f(2)]^2 f(3) = -3(-2) + [-2]^2 = 6+4 f(3) = 10 ---------------- f(4) = -3 f(3) + [f(3)]^2 f(4) = -3(10) + 10^2 = 100-30 etc ....